# JUREX-4E: Juridical Expert-Annotated Four-Element Knowledge Base for Legal Reasoning

### v1

@@ 4 Data Distribution
SB 23.12 27 51.64 17
OB 15.86 15 36.01 25
SA 28.00 30 42.38 21
OA 48.45 45 342.5 230
Table 1: Comparison of Legal Element Lengths: LLM vs. Expert. SB = Subject, OB = Object, SA = Subjective Aspect, OA = Objective Aspect.
As shown in Table 1, we compare the length of legal elements between expert-annotated descriptions in JUREX-4E and LLM-generated outputs across 105 charges that overlap with the Lecard-V2 dataset Li et al. (2024c), which is one of the most comprehensive legal datasets, covering 184 criminal charges. We find that:
(1) The average total length of expert annotations (472.53) is more than four times longer than that of LLM-generated outputs (115.43), indicating that the former include more detailed information.
(2) The median difference between the Subject (SB), Object (OB), and Subjective Aspect (SA) is relatively small, as these elements are typically fixed. For example, the SB is often a general entity, and the SA is often intent or negligence.
(3) The median and mean values for SB and SA in the expert annotations differ, especially for SB (17 v.s. 51.64). This discrepancy arises because certain specialized charges may require more detailed explanations. For example, in the crime of copyright infringement, the definition of “work” under the subject element has 9 occasions. Detailed data distribution for each element is provided in Appendix A.
(4) The main difference between Expert and LLM is in the Objective Aspect (342.5 v.s. 48.45 in Mean). This is because the OA includes a range of factual elements describing the criminal behavior, such as the conduct, object, result, time, and location, which are most emphasized in legal provisions and are central to various legal interpretive theories.
@@ 5 Human Evaluation
We selected 6 complicated crimes in Chinese judicial practiceOuyang et al. (1999) to evaluate whether the LLM can handle the Four-element Theory. Drawing from previous workDeng et al. (2023); Cui et al. (2024); Zhou et al. (2023), we define LLM-generated knowledge as information produced by the LLM based on its pre-trained knowledge and contextual prompts. For detail, we provide the LLM with legal articles and the definition of each element in FET, prompting it to generate the four-elements base on these metrical. The LLM is expected to autonomously identify and generate the four elements based on its learned understanding of legal concepts.
We invite legal experts to assess the four elements generated by the LLM from four dimensions: Precision, Completeness, Representativeness, and Standardization.
  • Precision: Whether the key components of each element are accurately identified. This dimension mainly evaluates whether the four elements faithfully represent the legal provisions.
  • Completeness: Whether all necessary information of each element is included. This assesses whether any essential content is missing, such as the omission of a description for specific subjects, like government officials.
  • Representativeness: Whether the annotations highlight the most critical scenarios in judicial practice. For example, in crimes of intentional injury, this would involve describing the representative means of harm.
  • Standardization: Whether the four elements are clearly defined, ensuring consistency in the expression of identical elements across different crimes (e.g., consistent description of general subjects), with concise and easily understandable explanations, free from legal ambiguities or misunderstandings.
Each dimension was scored by two types of experts: one group with a pure legal background and another group with a combined background in law and Artificial Intelligence, all of whom have passed the bar examination. The experts were selected to balance domain expertise and interdisciplinary perspectives. Scores were averaged across the two groups. Details about 1-5 scale criteria and annotator background are provided in Appendix C.
@@
Table 2: Performance comparison of four elements across methods. $\delta$ represents the score difference between expert and LLM-generated four-elements, with experts outperforming LLMs in all dimensions.
As shown in Table 2, expert annotations consistently outperform LLM-generated elements across all four dimensions, highlighting the limitations of LLMs in understanding legal elements. The most pronounced deficiencies are observed in Completeness (+0.86) and Representativeness (+0.88). This suggests that while LLMs can generate formally standardized and relatively accurate four elements, their description are not specific enough and do not adequately reflect the representative features of a charge’s criminal composition.
6 Evaluate Expert Knowledge on Charge Disambiguation
In the preceding section, the human evaluation demonstrated that experts annotated higher-quality four-elements. To further quantitatively evaluate the annotations, a direct way is to judge whether different charges can be distinguished according to the four-element definition of crime constitution. Therefore, we introduce the Similar Charge Disambiguation (SCD) taskYuan et al. (2024); Li et al. (2024a).
6.1 Experiment Settings
6.1.1 Dataset
We chose the dataset released by Liu et al. (2021), which includes five charge sets with the largest number of cases. To evaluate performance on representative tasks, we selected three 2-label classification groups commonly examined in other datasets Yuan et al. (2024): Fraud & Extortion (F&E), Embezzlement & Misappropriation of Public Funds (E&MPF), and Abuse of Power & Dereliction of Duty (AP&DD). Each crime has over 1.9k cases, with a total of 13,962 cases. The details of the classification groups are shown in Appendix D. Following previous work Liu et al. (2021); Yuan et al. (2024), we use Average Accuracy (Acc) and macro-F1 (F1) as evaluation metrics.
6.1.2 Baselines and Methods
To evaluate SCD tasks, we consider two ways of incorporating legal knowledge. The first directly integrates legal statutes, represented by GPT-4o Achiam et al. (2023) as the baseline and GPT-4o+Article, which explicitly provides relevant legal articles to the model. The second adopts structured legal reasoning to enhance interpretability and accuracy. We consider Legal-CoT, a Chain-of-Thought Kojima et al. (2022) variant that conducts a stepwise analysis based on the FET, and MALR Yuan et al. (2024), a multi-agent framework that decomposes legal tasks into sub-tasks in four-element structures. Details of each baseline are provided in Appendix 7.
We use an unified approach to introduce four-element descriptions. For each group of similar charges, the model receives charges’ four-elements from JUREX-4E or generated by LLM to aid classification. Specifically, GPT-4o+FETExpert relies on expert-annotated four-elements, while GPT-4o+FETLLM relies on LLM-generated four-elements. As shown in Appendix D, the instruction format is consistent across methods, with only the [Four Elements of candidate charges] varying based on the source. All experiments are conducted in a zero-shot setting, with the max_tokens set to 3,000 (or 10,000 for COT and MALR reasoning) and temperature set to 0 or 0.0001(In repeated experiments).
6.2 Results
Model F&E E&MPF AP&DD Average
@@ 6.2 Results
GPT-4o+FETLLM 95.73 96.56 91.87 92.01 89.61 89.69 92.40 92.75
@@
Table 3: Results of Charge Disambiguation. FET means introducing the Four-element theory with knowledge obtained from experts and LLM method. Highest results are in bold.
As shown in Table 3, the GPT-4o+FETExpert performs best in discriminating similar charges, indicating that expert annotation is superior to other methods of directly or indirectly introducing FET with LLMs. Specifically, we can derive the following observation:
Effectiveness of Domain-Specific Legal Knowledge:
Among all approaches, those that explicitly incorporate domain-specific legal knowledge, such as GPT-4o+Article, Legal-CoT, and MALR, outperform GPT-4o alone. This highlights the importance of integrating legal knowledge.
Importance of Concrete Four-element Knowledge:
The accuracy of both Legal-CoT and MALR is still lower than GPT-4o+FET methods. This suggests that, compared to embedding the Four-Element Theory into LLMs’ reasoning process, providing concrete charge four-elements enables the model to better understand the different crimes’ composition.
Superiority of Expert Annotations:
Compared with the indirect introduction of FET reasoning, the method of directly introducing four-elements to the model (GPT-4o+FET) achieves better results. Notably, GPT-4o+FETExpert surpassing the GPT-4o+FETLLM by 0.65 in average accuracy and 0.70 in average F1-score, underscoring the superior quality and reliability of expert annotations in legal tasks, aligning with human evaluations in Table 2 and reaffirming the critical role of human expertise in legal decision-making.
7 Can Expert Knowledge Benefit More Downstream Tasks?
In this section, we design a simple framework to apply the expert-annotated four elements to Legal Case Retrieval (LCR), a task in which relevant cases are retrieved based on given facts. It is an important step in the practice of analyzing cases and making judgments, and it requires the precise application of the four-element theory to matches cases with similar criminal compositions.
Figure 2: An expert-guided FET method to enhance legal case retrieval by incorporating expert four elements.
7.1 Method
@@
BERT 0.1511 0.1794 0.1978 0.0199 0.0753 0.1299 0.2157 0.2579 0.1136
ChatLaw 0.2049 0.2328 0.2745 0.0353 0.1306 0.1913 0.2684 0.3751 0.1285
GEAR * * * 0.0630 0.1706 0.3142 0.4625 * 0.2162
BGE 0.4737 0.5539 0.5937 0.0793 0.2945 0.4298 0.6500 0.7394 0.1926
FETLLM 0.5139 0.5862 0.6291 0.0980 0.2967 0.4769 0.6802 0.7828 0.2140
        - base 0.3583 0.4293 0.4798 0.0506 0.2240 0.3644 0.5383 0.6652 0.1453
FETExpert_guided 0.5211 0.5920 0.6379 0.1024 0.3049 0.4883 0.6885 0.7967 0.2155
        - base 0.3766 0.4584 0.5111 0.0715 0.1894 0.3709 0.5891 0.7203 0.1624
Table 4: SCR results. Bold fonts indicate leading results in each setting. * denotes that the indicator is not applicable to the current model.
We implement a standard dense retrieval approach BGE using BGE-m3 Chen et al. (2023), an advanced embedding model for dense retrieval. Given a query $q$ and a candidate case $c$, their vector representations $\mathbf{v}_{q}$ and $\mathbf{v}_{c}$ are obtained through shared encoder $E$: $\mathbf{v}_{q}=E(q),\quad\mathbf{v}_{c}=E(c)$. We used the BGE-m3 model without fine-tuning as the shared encoder. Next, the relevance score is computed via cosine similarity:
$\text{sim}_{\text{base}}(q,c)=\frac{\mathbf{v}_{q}\cdot\mathbf{v}_{c}}{\|\mathbf{v}_{q}\|\|\mathbf{v}_{c}\|}$ (1)
To retrieve the top-k most similar cases, we rank the candidates based on their cosine similarity to the query. Denote the set of candidate cases as $C=\{c_{1},c_{2},\dots,c_{n}\}$, where $n$ is the total number of candidate cases. We compute the similarity for each $c_{i}\in C$, and select the top-k candidates with the highest similarity scores.
As shown in Figure 2, to leverages expert-annotated four elements of charges, we introduce an BGE+FETExpert_guided method for the retrieval process, consisting of three steps: (1) Predicting charges, a LLM $\mathcal{M}_{p}$ predicts potential charges $Z=\{z_{1},...,z_{k}\}$ from case facts. (2) Matching elements, retrieving corresponding charge’s four-elements $\{f_{z}\}_{z\in Z}$ in JUREX-4E. (3) Analyzing case facts. Guided by $\{f_{z}\}$, another LLM $\mathcal{M}_{g}$ generates case-specific four elements $a_{c}$ for candidate $c$. The final similarity score combines factual and theoretical alignment:
$\text{sim}_{\text{final}}(q,c)=\alpha\cdot\text{sim}_{\text{base}}(q,c)+(1-\alpha)\cdot\text{sim}_{\text{f}}(a_{q},a_{c})$ (2)
where $\alpha=0.7$ and $\text{sim}_{\text{f}}$ measures the similarity between the generated four-element descriptions.
To facilitate comparison, we also design a BGE+FETLLM method that directly prompt the LLM $\mathcal{M}_{g}$ with the concept of Four-Element Theory to generate case-specific four elements $a_{c}$.
7.2 Dataset
LeCaRDv2Li et al. (2024c) is the latest version of LeCaRDMa et al. (2021), which is widely used in LCR task Li et al. (2024b); Zhou et al. (2023). It comprises 800 queries and 55,192 candidates extracted from 4.3 million criminal case documents. There are two common evaluation settings for this dataset: one uses a subset Qin et al. (2024) with a candidate pool size of 1,390, while the other uses the full set Li et al. (2024c) with a candidate pool size of 55,000. We conducted experiments under both settings.
Following previous workFeng et al. (2024); Qin et al. (2024), we adopt commonly used evaluation metrics. For the subset, we use NDCG@10, 20, 30, Recall@1, 5, 10, 20, and MRR. For the full dataset, we use Recall@100, Recall@200, Recall@500, and Recall@1000.
7.3 Baselines
Consistent with earlier workLi et al. (2024c); Qin et al. (2024), we compare some dense retrieval methods, including: BERTDevlin (2018), LawformerXiao et al. (2021), ChatLaw-Text2Vec [^1] https://modelscope.cn/models/fengshan/ChatLaw-Text2VecCui et al. (2023), SAILERLi et al. (2023), GEARQin et al. (2024). Details of each baseline is shown in Appendix E. These baselines are implemented using the FlagEmbedding Toolkit [^2] https://github.com/FlagOpen/FlagEmbedding with a RTX 3090.

### v2

@@ 5.3 Data Statistics
Subject 23.12 27 51.64 17
Object 15.86 15 36.01 25
Subjective Aspect 28.00 30 42.38 21
Objective Aspect 48.45 45 342.5 230
Table 1: Comparison of element lengths.
Our dataset comprises 155 common criminal charges. These charges are selected based on their frequency in over 2.6 million publicly available criminal cases in China: specifically, we include all charges that appear more than 3,000 times, which together account for 91.71% of all cases (Appendix A), ensuring coverage of the most common real-world judicial scenarios.
To compare the quality of expert-annotated FETs (expert-FETs) and LLM-generated FETs, we selected 105 charges in JUREX-4E that also appear in the widely used LeCaRDv2 dataset Li et al. (2024d). LLM-generated FETs were produced using the same setup as before, with a maximum output of 8192 tokens. Table 1 summarizes the differences in element length, with full length distributions available in Appendix A.
Overall, expert-FETs are significantly longer, with an average total length of 472.53 words compared to 115.43 for LLM-generated ones. The most pronounced gap appears in the Objective Aspect (OA) (mean: 342.5 vs. 48.45), where experts provide detailed factual descriptions, such as action, result, time, and location, often underdeveloped in LLM outputs. While the Subject (SB), Object (OB), and Subjective Aspect (SA) show smaller median differences, notable variation remains, especially in SB (mean: 51.64 vs. 17), which in certain charges involves complex legal interpretations (e.g., “work” in copyright infringement) requiring more elaborate legal definitions.
@@ 6 Human Evaluation
To compare the quality of expert-annotated and LLM-generated FETs, we selected six complicated charges in Chinese judicial practice Ouyang et al. (1999). Based on prior theoretical framework Zhang (2007a), we assess the quality of FETs along four independent dimensions: Precision, Completeness, Representativeness, and Standardization.
  • Precision: Whether each element accurately aligns with its statutory definition, reflecting key terms in the corresponding legal article.
  • Completeness: Whether each element includes all practically necessary information, ensuring the definition is sufficient to guide legal reasoning.
  • Representativeness: Whether the annotations reflect the most typical and practically significant scenarios in judicial practice.
  • Standardization: Whether the expressions of elements are consistent across different charges, with clear, concise, and unambiguous language that facilitates understanding and minimizes interpretive variance.
Evaluation was conducted by experts from two backgrounds: one group with a purely legal background and another with a combined background in law and AI, all of whom have passed the National Judicial Examination. The experts were selected to balance domain expertise and interdisciplinary perspectives. Scores were averaged across the two groups. Details about the 1-5 scale criteria and annotator background are provided in Appendix E.
@@
Table 2: Performance comparison of four elements across methods. $\delta$ represents the score difference between expert and LLM-generated FETs, with experts outperforming LLMs in all dimensions.
As shown in Table 2, expert annotations consistently outperform LLM-generated elements across all four dimensions. The most pronounced deficiencies are observed in Completeness (+0.86) and Representativeness (+0.88). This aligns with our earlier analyses, where expert-generated elements include more factual details and representative descriptions. The gap in Precision (+0.57) suggests a tendency toward vague or legally irrelevant content, while the smaller difference in Standardization (+0.23) shows that LLMs can mimic structural patterns but lack deeper normative consistency. These results demonstrate the importance of expert supervision in providing reliable legal knowledge.
7 Evaluation on Similar Charge Disambiguation
To further validate annotation quality, we introduce the Similar Charge Disambiguation (SCD) task Yuan et al. (2024); Li et al. (2024a). Given the case fact and a set of similar charges, SCD task requires the model to identify which charge is correct. We evaluate whether similar charges can be effectively distinguished based on their FETs, and whether expert-annotated FETs perform better than LLM-generated FETs.
Model F&E E&MPF AP&$D$ Average
@@ 7 Evaluation on Similar Charge Disambiguation
Farui-plus+FET4o 89.09 90.27 86.32 88.00 75.90 77.67 83.77 85.31
Farui-plus+FETExpert 89.29 90.98 86.13 87.54 76.25 78.12 83.89 85.55
Qwen2.5-72b+FET4o 93.15 95.06 90.99 93.56 87.71 88.56 90.62 92.39
Qwen2.5-72b+FETExpert 93.29 95.18 91.18 93.66 87.81 89.45 90.76 92.76
GPT-4o+FETfarui 94.86 96.12 91.84 92.64 89.35 89.85 92.02 92.87
GPT-4o+FETqwen 95.53 96.53 91.82 92.96 89.48 90.09 92.28 93.19
GPT-4o+FET4o+farui+qwen 94.97 96.24 91.84 92.73 89.69 90.12 92.17 93.03
GPT-4o+FET4o 95.73 96.56 91.87 92.01 89.61 89.69 92.40 92.75
GPT-4o+FET4o + ICL 95.74 96.36 91.84 92.01 90.48 90.63 92.69 93.00
@@
Table 3: Performance on the Similar Charge Disambiguation (SCD) task. “Expert” refers to our expert-annotated FET, while “4o”, “qwen”, and “farui” refer to FET generated by different LLMs. Highest results are in bold.
7.1 Experiment Settings
7.1.1 Dataset and metrics
We chose the SCD dataset released by Liu et al. (2021). Following previous work Yuan et al. (2024), we selected three 2-label classification groups: Fraud & Extortion (F&E), Embezzlement & Misappropriation of Public Funds (E&MPF), and Abuse of Power & Dereliction of Duty (AP&DD). Each charge has over 1.9k cases, with a total of 13,962 cases. The details of the groups are shown in Appendix F. Following previous work Liu et al. (2021); Yuan et al. (2024), we use Average Accuracy (Acc) and macro-F1 (F1) as evaluation metrics.
7.1.2 Baselines and Methods
We compared the following baselines: GPT-4o Achiam et al. (2023) and GPT-4o+Article, which explicitly supplies relevant legal articles; Legal-COT Kojima et al. (2022), a Chain-of-Thought variant that applies the Four-Element Theory step by step, and MALR Yuan et al. (2024), a multi-agent framework that decomposes legal tasks into FET-aligned subtasks. Details are in Appendix 9.
Methods: Following Section 4, our main model is GPT-4o. We also compared Farui-plus (the latest version of Tongyifarui, representative legal LLM) and Qwen2.5-72B Bai et al. (2023) (representative open-source LLM). To incorporate FET knowledge, each group of similar charges is augmented with four-element descriptions, either generated by LLMs or sourced from JUREX-4E. For example, GPT-4o+FETLLM uses LLM-generated FETs, while GPT-4o+FETExpert uses expert-annotated ones. The input format is fixed across methods, differing only in the [four-elements of candidate charges] (Appendix F). All experiments are zero-shot, with max_tokens set to 3,000 (10,000 for Legal-COT and MALR) and a temperature of 0 or 0.0001 in repeated runs.
To further explore generation setups, we also evaluate an ICL variant, GPT-4o+FET4o + ICL, where two representative FET exemplars (Theft and Snatching) are provided in the prompt to guide generation.
7.2 Results
The SCD results are shown in Table 3, where we can observe that:
Effectiveness of Structured FET Knowledge: Providing specific structured charge FETs yields the highest accuracy among all legal knowledge integration methods. Compared to implicit approaches, such as prompts (GPT-4o+Article, Acc 92.09) or reasoning chains (Legal-COT, Acc 89.95), structured FET knowledge offers more effective support for legal decision-making (e.g., GPT-4o+FETExpert, Acc 93.05)
Superiority of Expert-Annotated FET: Expert-annotated FET consistently outperforms LLM-generated FET across three representative LLMs, including FETfarui, FETqwen, FET4o, and their combination (FET4o+farui+qwen). For example, GPT-4o+FETExpert surpasses GPT-4o+FET4o by 0.65 in average accuracy and 0.70 in F1-score.
Consistent Gains Across Models: Expert-annotated FETs yield consistent performance gains across different SCD models. When applied to Farui-plus, Qwen2.5-72b, and GPT-4o, it improves F1-score by +0.24, +0.37, and +0.70, respectively over their LLM-generated FET baselines.
The ICL-based variant yields consistent improvements over direct prompting (Table 3), demonstrating the benefit of exemplar guidance, though it still lags behind expert-FETs. We also conducted McNemar’s test on paired samples between each LLM-generated FET and the expert-FET, which shows statistically significant improvements ($p<0.05$) across all tasks (see Table 11 in Appendix F).
8 Application in Legal Case Retrieval
In this section, we design a simple expert-guided FET method to apply JUREX-4E to the Legal Case Retrieval (LCR) task, which retrieves relevant cases based on case facts. This task is well-suited for FET because it requires a comprehensive comparison of the four elements across different charges in cases.
@@ 8 Application in Legal Case Retrieval
BGE (case_fact only) 0.4737 0.5539 0.5937 0.0793 0.2945 0.4298 0.6500 0.7394 0.1926
BGE+FET (Qwen2.5) 0.5125 0.5858 0.6350 0.1104 0.2870 0.4653 0.6679 0.7836 0.2168
   FET only 0.3367 0.3971 0.4487 0.0622 0.2006 0.3279 0.4806 0.6037 0.1524
BGE+FET (Expert, Qwen2.5) 0.5295 0.5979 0.6416 0.1124 0.3122 0.4838 0.6791 0.7824 0.2206
   FET only 0.3354 0.4035 0.4541 0.0849 0.1923 0.3076 0.4839 0.6097 0.1606
BGE+FET (GPT-4o) 0.5139 0.5862 0.6291 0.0980 0.2967 0.4769 0.6802 0.7828 0.2140
   FET only 0.3583 0.4293 0.4798 0.0506 0.2240 0.3644 0.5383 0.6652 0.1453
BGE+FET (Expert, GPT-4o) 0.5211 0.5920 0.6379 0.1024 0.3049 0.4883 0.6885 0.7967 0.2155
   FET only 0.3766 0.4584 0.5111 0.0715 0.1894 0.3709 0.5891 0.7203 0.1624
Table 4: Performance on the Legal Case Retrieval (LCR) task. The highest results are in bold. “FET only” indicates using the four-element descriptions without case facts.
8.1 Dataset and Metrics
LeCaRDv2 Li et al. (2024d) is the latest version of LeCaRD Ma et al. (2021), which is widely used in legal tasks Li et al. (2024b); Zhou et al. (2023). It comprises 800 queries and 55,192 candidates extracted from 4.3 million criminal case documents. Following previous work Qin et al. (2024), we chose 1390 candidates and used NDCG@10, 20, 30, Recall@1, 5, 10, 20, and MRR as metrics. We also tested different candidate pool settings (see Appendix G). The results are consistent.
8.2 Baselines and Methods
We adopt a dense retrieval framework based on BGE-m3 Chen et al. (2023), a strong embedding model for legal and general-domain texts. Given a query $q$ and a candidate case $c$, we compute their vector representations using a shared BGE-m3 encoder. Retrieval is performed by computing cosine similarities between the query and all candidates and selecting the top-k candidates.
To enhance retrieval accuracy, we compare the following three methods:
(1) BGE(case_fact only): Standard dense retrieval using only BGE-m3 embeddings of the raw case facts.
(2) BGE+FET ($\mathcal{M}_{g}$): We prompt different LLMs $\mathcal{M}_{g}$ to generate a structured four-element description of each case (case-FET) based solely on its facts, without using external knowledge. These case-FETs are then embedded with BGE-m3 and used to compute similarity. Because the FET abstracts away case-specific details, we combine the original fact-based similarity and the FET-based similarity in a ratio of 7:3.
(3) BGE+FET (Expert, $\mathcal{M}_{g}$): An expert-guided FET method that incorporates JUREX-4E to guide case-FET generation. It consists of four steps:
  1. 1.
    Charge Prediction. A charge prediction model $\mathcal{M}_{p}$ (Qwen-plus, details see Appendix D) predicts the set of likely charges $Z=\{z_{1},...,z_{k}\}$ for the query case.
  2. 2.
    Expert-FET Matching. Retrieving corresponding charge’s four-elements $\{f_{z}\}_{z\in Z}$ for each predicted charge in JUREX-4E. These provide theoretical guidance for subsequent reasoning.
  3. 3.
    Case-FET Generation. Guided by $\{f_{z}\}$, the LLM $\mathcal{M}_{g}$ generates case-specific four-elements $fet_{c}$ for candidate $c$.
  4. 4.
    Dense retrieval. We embed the generated FETs using BGE-m3 and compute similarity scores as in Method (2), combining both factual and FET-based similarities.
For the $\mathcal{M}_{g}$, we chose Qwen2.5-72b and GPT-4o. The retrieval framework is implemented with the FlagEmbedding Toolkit [^4] https://github.com/FlagOpen/FlagEmbedding with an RTX 3090. Following prior work Li et al. (2024d); Qin et al. (2024), we also compare some dense retrieval methods to examine the representativeness of BGE-m3. Results of baselines and prompt templates are available in Appendix G.